Statistics Unit Three Let z be a random variable with a standard normal distribution. Find the indicated probability (Round your answer to four decimal places.) Pz z 1.21)- Shade the corresponding area under the standard normal curve “3-2 or 032 -3-2 ??
Answer #1

We have to find P(Z \geq 1.21)

The first graph is correct. Greater than means right side area.

P(Z \geq 1.21) = 1 - P(Z < 1.21) = 1 - 0.8869 = 0.1131